OpenStudy (anonymous):

y= ln(e^(x^2-1) find y'(1) a. 0 b. 1/2 c. 1 d. 2 e. undefined PLEASE SHOW ALL STEPS THANK YOU!

OpenStudy (anonymous):

(d)

OpenStudy (anonymous):

\[\ln(e^{x^2-1}) = x^2-1\] so \[y=x^2-1\] and \[y'=2x\] so y'(1) = 2.

OpenStudy (anonymous):

ln(e^(x^2-1))=x^2-1, differentiate this and get dy/dx=2x, so at 1 dy/dx=2